# Vertex Block Descent

### v2

@@ 4.5. Comparisons to XPBD
XPBD replaces the Hessian matrix and uses only the Hessian of inertia potential. This omission is justified by using a small time step, because the significance of the inertia potential increases quadratically as the time step decreases. Nonetheless, with complex examples, the impact of this approximation can be severe, even with small time step. This is demonstrated in Figure 18 with a challenging collision-rich scenario involving a squishy ball with tentacles dropped to the ground. Comparing XPBD with 120 iterations per step (17(a)) to our method with the same number of iterations (17(d)), we can see that our method not only achieves a more stable animation, it also performs faster because of its improved parallelism, as compared to XPBD. Reducing the time step helps XPBD even when using a similar total number of iterations per frame (17(b)). However, simply reducing the time step is not sufficient in this case, as XPBD also needs more frequent collision detection (17(c)). Using collision detection with the same frequency as ours while taking small time step (17(b)) leads to collisions that are detected too late and cause stability issues in this case. This is not only because the collisions that are detected too late are deeper, but also because smaller time step lead to higher vertex velocities when resolving stiff collisions.
@@ 5.3. Unified Simulations
6. Limitations
@@
VBD is a primal solver (Macklin et al., 2020), so it can easily handle high mass ratios (see Figure 6, 19, and 20), but it struggles with high stiffness ratios. This is shown in Figure 23 using a stiffness ratio of 1:10000, where VBD has poor convergence behavior.

### v4

@@ 5.5. Comparisons to XPBD
XPBD replaces the Hessian matrix and uses only the Hessian of inertia potential. This omission is justified by using a small time step, because the significance of the inertia potential increases quadratically as the time step decreases. Nonetheless, with complex examples, the impact of this approximation can be severe, even with small time step. This is demonstrated in Figure 19 with a challenging collision-rich scenario involving a squishy ball with tentacles dropped to the ground. Comparing XPBD with 120 iterations per step (19(a)) to our method with the same number of iterations (19(d)), we can see that our method not only achieves a more stable animation, it also performs faster because of its improved parallelism, as compared to XPBD. Reducing the time step helps XPBD even when using a similar total number of iterations per frame (19(b)). However, simply reducing the time step is not sufficient in this case, as XPBD also needs more frequent collision detection (19(c)). Using collision detection with the same frequency as ours while taking small time step (19(b)) leads to collisions that are detected too late and cause stability issues in this case. This is not only because the collisions that are detected too late are deeper, but also because smaller time step lead to higher vertex velocities when resolving stiff collisions.
@@ 6.3. Unified Simulations
7. Discussion
We derive our method as a block coordinate descent method for variational time integrators, which offers optimization techniques like PSD projection and line search. The fact that we do not require those techniques to guarantee stability actually makes our method a more general solver of nonlinear equations. When line search is not used, our method can effectively manage non-conservative forces, such as friction, the same as how it handles conservative forces. In other words, our method allows for a seamless transition between block coordinate descent and block Gauss-Sediel (Hageman and Porsching, 1975; Grippo and Sciandrone, 2000). While we do not practically utilize these optimization techniques derived from the descent view, they remain available options for users.
@@ 7. Discussion
VBD is a primal solver (Macklin et al., 2020), so it can easily handle high mass ratios (see Figure 6, 20, and 21), but it struggles with high stiffness ratios. This is shown in Figure 24 using a stiffness ratio of 1:10000, where VBD has poor convergence behavior.
@@ 8. Conclusion
Acknowledgements.
We thank Miles Macklin and Theodore Kim for an inspiring discussion on primal solvers. This work was supported in part by NSF grant 1764071.